A can paint a certain house in 10 days and B can paint the same house in 12 days. How long will it take to paint the house if both men work?: A can paint a certain house in 10 days and B can paint the same house in 12 days. How long will it take to paint the house if both men work?
Answer by Alan3354(1415)

A can paint a certain house in 10 days and B can paint the same house in 12 days. How long will it take to paint the house if both men work?
---------------------
A can paint 1 house in 10 days, or 1/10 of the house per day.
B can paint 1/12 of the house per day.
So, working together, the paint 1/10 + 1/12 per day.
1/10 + 1/12 = 11/60 of the house per day.
Days/house = 60/11 = 5 5/11, or apx 5.4545....
---------------------------------------
This looks good, because it's less than either of them working alone. If they had both taken 10 days, the answer would be 5. If both 12 days, the answer would be 6, so the answer should be between 5 and 6, which it is.